# Prompts Are Programs Too! Understanding How Developers Build Software Containing Prompts

### v1

@@ 3.2. Participants
A summary of our participants is in Table 1. The inclusion criterion was developers who had created at least one prompt program before, according to the definition presented in Section 1. Our participants were men () and women () from a broad range of roles in technology, including Machine Learning Engineer, Senior Software Engineer, Ph.D. Student, and Principal Data and Applied Scientist. All participants had prior programming experience (median 8.5 years) and had written a prompt program before (median 10 programs). Participants also regularly used foundation models: 10 participants reported using them more than once daily, 5 participants reported using them once daily, and 5 participants reported using them weekly.
@@ 3.3. Interview Protocol
The interviews were 60 minutes long and were conducted over Zoom. The procedure was approved by our Institution’s Review Board. Participants were compensated with a $20 Amazon gift certificate.
To begin, the participant completed a demographic and background survey with questions such as gender, number of years of programming experience, and number of prompt programs developed. We follow best practices in reporting gender in HCI (Scheuerman et al., 2020). Then, the interviewer presented the participant with the definition of a prompt program and asked the participant to recall the most recent time they wrote a prompt program. If the participant had access to the prompt, they retrieved the prompt and any associated history with the prompt. The participant then provided a brief overview of the prompt. To re-familiarize the participant with their prompt and their process, the participant discussed the prompt’s design choices (i.e., a decision made in how to implement the prompt). Next, they discussed the challenges they faced while developing the prompt. Finally, the participant discussed the overall process they used to develop the prompt.
To develop the interview protocol, we identified several themes to explore: data, requirements, design, implementation, evaluation, debugging, and deployment based on the broad literature review and the authors’ prior experience developing prompt programs (see Section 3.1). We then generated questions focused on each theme in the interview protocol. The challenges were discussed before the process so it was not biased by the discussion about the prompting process. A sample of the interview questions are included in Figure 2; the full interview protocol and the demographic and background survey are included in the supplemental materials (Authors, 2024).
Following best practices in software engineering research (Ko et al., 2015), we piloted the interview protocol with three individuals who had developed prompt programs to clarify the wording of the survey and interview as well as validate our definition of prompt program.
@@ 3.4. Threats to Validity
Participants could have memory biases that could introduce errors in their recounting of their prompt design choices and development experience. To reduce this threat, we asked participants to recount their most recent prompt that met the definition of a prompt program. When possible, we asked participants to review their prompt to ground their answers. Further, the authors’ own previous experiences with prompt development could also introduce confirmation biases of the developed theory. Overall, we reduced the limitations of this study by performing triangulation with prior literature and validating the theory with study participants and a professional who have engaged in prompt programming.
@@
Using snowball sampling and recruiting within the authors’ social networks may introduce sampling bias, so our sample may not be representative of all developers creating prompt programs. Additionally, self-selection bias could influence the results. Interviews were conducted in English and could cause less representation from non-native English speakers. Also, some participants were unable to disclose all details of their process due to company policy. Thus, the results from this study may not fully generalize beyond our sample of participants.
@@ 4. Properties of Prompt Programming
We find that prompt programming is the interaction between three entities: the developer, the foundation model, and the prompt. To write a prompt, a programmer uses their prior experience with the FM to construct a mental model of how the FM might perform on the task. This mental model informs how the prompt is written. After the prompt is run, the developer observes the generated output to create a hypothesis or new belief about the model’s behavior, which then is used to update the developer’s mental model. The updated mental model is then used to influence the next prompt version.
Thus, prompt programs are both very flexible in terms of accepting a range of inputs, but also very overfit as it is the product of the specific FM being used, task at hand, and observations of the developer. We first discuss the aforementioned properties of prompt programming in further detail (see Section 4.1) and end with a discussion of its implications (see Section 4.2).
4.1. Results
4.1.1. Programmer
@@
Programmers develop a mental model about the FM’s behavior on the prompt to predict how it may behave.
Technology users form mental representations of their interactions with devices, which can be developed through hands-on experience with the technology (Staggers and Norcio, 1993). Understanding the ML model’s abilities is important but challenging for developing ML-enabled systems (Giray, 2021; Wolf and Paine, 2020; Dolata and Crowston, 2023; Arpteg et al., 2018), including with FMs (Arawjo et al., 2023; Jiang et al., 2022; Wu et al., 2022a; Hassan et al., 2024). Just as developers form mental models of code by exploring the codebase or talking to colleagues (LaToza et al., 2006), participants described developing a mental model of the FM’s behavior on the prompt by running the prompt. All participants developed their mental model by observing the FM’s performance, either by examining individual examples (P6, P7, P8, P9, P10, P12, P14, P15, P16, P17, P18, P19, P20) or looking at metrics (P4, P5, P7, P8, P9, P11, P14, P15, P16, P18, P19). Based on these observations, the participant formed a hypothesis or belief about the model’s behavior and updated the prompt accordingly. This often included creating a new explicit instruction, “guideline” (P8), “rule” (P14), or “mandate” (P20) in the prompt to address the observation (see Section 5.1.3), as noted by Zamfirescu-Pereira et al. (2023b).
The programmer’s mental model was the accumulated beliefs about the FM’s behavior. These beliefs could differ or even contradict between individuals. For instance, while some participants (P1, P4, P5, P6, P9, P15, P16, P18, P20) included examples in the prompt for few-shot learning (Brown, 2020) to achieve better results, others (P7, P13, P19) believed the inclusion of examples caused the prompt to overfit to those examples: “[we did not] use few-shot prompting because…it over-indexes on the examples. After a while, you start seeing very repetitive and monotonous output” (P7). This overfitting of examples has been found in NLP literature (Liu et al., 2023) and has also been noted in practice (Jiang et al., 2022).
Programmers’ mental models are not reliable.
Mental models are known to be incomplete and inaccurate (Staggers and Norcio, 1993). Due to the lack of transparency of neural models (Arpteg et al., 2018) like FMs in prompt engineering (Mishra et al., 2023), participants mentioned they were not always confident about their developed mental models (P2, P4, P5, P6, P8, P9, P10, P15, P17, P18, P19): “It’s a black box model. Nobody knows what’s going on inside. It’s a science of faith” (P17). Additionally, the stochastic nature of FMs (Bender et al., 2021) made it difficult for participants to predict the model’s behavior, as found in prior work (Jiang et al., 2022): “I have never been confident [in predicting an FM’s behavior]…and I don’t think I ever will” (P10).
@@
Programmers use external knowledge sources and prior experience with FMs to develop their mental model.
Prior to prompt programming, participants (P1, P2, P4, P11, P12, P15, P16, P18) developed their mental model of the FM through “prompt intuition” (P15) (i.e., prior experience of interacting with FMs): “I knew what works generally just being in the field” (P8). Literature also points to external knowledge sources as helpful for prompt programming (Parnin et al., 2023; Jiang et al., 2022). Participants used official resources, such as guides from the organization or AI companies like OpenAI (P6, P10, P14, P15), online courses (P7), research papers (P1, P2, P5); informal resources, like blogs, articles (P5, P10, P16, P20), and Reddit or Twitter posts (P5, P16, P20) written by “everyday users” (P20); expert colleagues (P1, P14, P15); and other example prompts (P16, P18).
While these resources were helpful, participants (P2, P10, P14, P15, P16, P20) expressed doubt as to whether they improved the prompt: “I don’t know if the best practices are the best. I know they are there for a reason, but sometimes they just don’t actually make that much of a change. There’s a reason the [expert] is saying that… I just take it they know better than me” (P14). Even company-specific resources did not help: “Even with following these best practices…put out internally by the company, some of them just straight up didn’t work” (P10). As a result, some participants felt like there were no such thing as best practices: “There are no standards for what’s a good prompt or what’s a bad prompt. It all seems like a haphazard science of just adding a word and maybe it will look better” (P17).
4.1.2. Foundation Model
@@
Participants ran their prompts on models that varied in capabilities (see Table 1). Some qualities were explicit or obvious; this included the types of input the model accepted (e.g., images and text (P9)), output generated (e.g., “GPT-3.5 model’s JSON output [mode]” (P7)), context window size (P2, P3, P8, P19), prompt formatting requirements (P6), generation speed (P4, P11, P15, P19), and the privacy of the model (P4).
@@
Some capabilities were latent and were discovered by interacting with the model, as noted in prior work (Arawjo et al., 2023; Jiang et al., 2022). Literature suggests FMs as having latent capabilities, as GPT’s performance has improved on some tasks and regressed on others over time (Chen et al., 2023; Ma et al., 2024). Participants found some differences between FMs to be qualitative: “[One] model was capturing some instructions in the prompt in a much more focused manner, whereas the [other] was a little more lax” (P15). In our sample, participants (P1, P4) noted a quality difference in open-source models (e.g., Llama (Touvron et al., 2023)) compared to closed-source models (e.g., GPT-4 (Achiam et al., 2023)): “Getting [the prompt] to work on a smaller model was very frustrating when with a zero-shot prompt, GPT-4 would almost achieve a hundred percent performance” (P4). Participants (P1, P9) also noted some FMs were unable to complete the task: “[The model was] generating garbage, and [after 30 iterations] I concluded the model is trash” (P1).
4.1.3. Prompt
@@
Prompts are sensitive to small details. Additionally, since prompts are a product of the programmer’s observations about model behavior as well as the qualities and capabilities of the model, we observe that prompts can be fragile.
Minute details in the prompt matter.
@@
Participants reported developing several techniques to influence the FM into generating the desired output. In addition to using known, high-level techniques from literature (e.g., assigning personas, clearly describing the task, selecting a prompting strategy, and providing data context) (ope, 2024c; Lin, 2024), participants described a range of subtle strategies they employed to obtain the desired output. This included specific phrasing for clarity or generality (P5, P13, P14, P16, P20); repeating phrases for emphasis (P4, P6, P10, P20); avoiding negations (P2, P12); or adding specific characters, like emojis, to encourage their appearance (P6, P10). Other participants noted strategies to visually format the prompt for organization purposes. This included using numbered or bulleted lists (P8, P9, P10, P16) as well as new lines and spacing (P1, P3, P5, P11, P12, P18). Other participants used techniques to emphasize specific parts of the prompt, including bolding (P10, P18) and capitalization (P7, P10, P12, P20).
@@
Participants (P10, P12) also converged on creative details to include in the prompt for better performance. For instance, to prevent the FM from leaking prompts, one participant “bullied GPT” (P10): “We have a very light threat in there saying, ‘If you share any of this information we’ve given it, you will get in a lot of trouble, I will get in a lot of trouble and it will cause great career harm.’ That worked for the most part” (P10). Another participant summarizing code told the FM to explain the code “‘so a sixth grader who doesn’t understand programming can understand it’. The explanations…were verbose, so I told it the description should be short…It was still super verbose. Then I added, ‘sixth graders don’t really like reading’ and for some reason, that got it to be shorter” (P12).
@@
Finally, participants noted the input and output data format impacted the performance of the prompt. Consistent with prior studies (Parnin et al., 2023; Liu et al., 2024), participants selected data formats that were structured and standardized, such as JSON (P3, P7, P9, P10, P11, P18), Markdown (P2, P7, P12), and XML tags (P1, P3, P7, P10, P20). Participants also noted instances where they would change the data representation to achieve better performance (P4, P13, P20): “[What worked better for us is] describing the APIs as though they are Python functions and having the documentation in a similar format to publicly available Python functions” (P13).
Prompts are finicky and fragile.
@@
As noted in NLP literature (Jiang et al., 2020; Ye et al., 2024) and empirical studies on prompting (Liu et al., 2023; Zamfirescu-Pereira et al., 2023b), a prompt’s performance varies based on its wording. In practice, participants found prompts to be “finicky” (P4), “fragile” (P20), “sensitive” (P1, P13), and “temperamental” (P15), since prompts would suddenly worsen in performance during prompt development, as corroborated in prior studies on prompt engineering (Parnin et al., 2023; Zamfirescu-Pereira et al., 2023b, a; Jiang et al., 2022; Wu et al., 2022a; Hassan et al., 2024). This could be due to a model’s ability to follow instructions, which can change over time (Chen et al., 2023). For instance, participants noted performance regressions in the prompt (P2, P8, P10, P12, P15, P16, P19, P20), most often by adding in “a bunch of new requirements” (P2) that would cause previous instructions to suddenly “deactivate” (P20) or be “disregarded” (P8) in the new version: “if I provide multiple guidelines, [the FM] would follow some and disregard others, especially when they’re conflicting or overlapping in nature” (P8).
@@
While prior NLP studies point to prompts being transferrable between FMs (Liu et al., 2023), participants found prompts to be “finicky between models” (P4), causing the prompt to break on the new FM and requiring additional development effort (P4, P10, P15, P18): “At one point, we had to do…a minor model version change… That broke almost everything. We had to rewrite so many prompts” (P10). This is because a change in the model could result in “a different output…or format. So that affects my whole flow. I have a regex converting [the output]…and the whole thing gets broken” (P18). This has also been found in prior prompt engineering work (Dolata et al., 2024).
Finally, chaining prompts was challenging due to the finicky nature of prompts, which has been observed in other studies (Wu et al., 2022a; Dolata et al., 2024): “Two changes that do really well separately could merge and become a terrible experience for unexplainable reasons” (P13). Participants (P8, P12, P13, P19) found that integration often resulted in breakage and required significant updates to the prompts so they could work with each other. When working individually, P19 focused on performing regular integration testing of multiple prompts: “[Integration testing is] highly correlated with how it’s going to perform in a production environment. In the end, we are going to integrate everything” (P19). However, collaboration could pose additional challenges in such scenarios, since programmers may be unable to perform tests with other concurrent in-progress prompts during development. P13 reported changing how she would “split up the work” (P13) with colleagues and feel an increased “rush to get my PR in before the other one does” (P13) to avoid being stuck with prompt integration.
4.2. Implications
Are the “best practices” best practices?
In our study, participants’ mental models on FMs conflicted, such as whether few-shot learning (Brown, 2020) was beneficial to prompt performance. While including examples is listed within the official prompting guide made by OpenAI (ope, 2024c), some of our participants (P7, P13, P19) were skeptical of this practice after observing the FM. This underscores the importance of the observations of developers in prompt programming, especially in the face of evolving model behaviors (Chen et al., 2023; Ma et al., 2024). However, such knowledge and observations are rarely shared, are siloed in informal social media posts (Parnin et al., 2023), and are not systematically collected and centralized. However, these observations provide valuable information about FM behavior, which at scale, could reveal the capabilities of FMs overall. Future tools could collect these observations and reasons for prompt updates. This could provide rich information for prompt provenance and version control for the individual programmer, as well as produce generalizable knowledge about the FM’s behavior at scale that could be used for automatic prompt program repair.
Rethinking prompt reuse
Due to the fragility of prompts and the unique qualities of each FM, reusing prompts without modification is difficult. Prompt reuse at higher levels of abstraction, such as template-based approaches (White et al., 2023, 2024), could be an effective way for prompt reuse. Community-based solutions for prompt sharing, like ShareGPT (sha, 2024), LangSmith Hub (lan, 2024), and Wordflow (Wang et al., 2024), could be a promising avenue for prompt reuse. These tools should facilitate effective search and retrieval of prompt programs, which to the best of our knowledge has not been studied. Future work could examine how prompt programmers query for previous prompts or outputs for reuse.
@@ 4.2. Implications
| Requirements | • Expressed outside of code | • Expressed in a prompt, which is also the implementation |
| • Dependent on the performance of the FM on the task | ||
| Design | • Base component: Class/function | • Base component: Single prompt |
| • Chained prompts are tightly coupled | ||
| Implementation | • Write a program that implements a specification in code | • Write a prompt that is a natural language specification that when ran on a foundation model, implements the specification |
| • Focus on rapid experimentation | ||
| Debugging | • Fault can be isolated to a single line of code | • Fault localization is not certain |
| • Can use debugging tools for more systematic debugging | • No debugging tools; reliance on shotgun debugging | |
| Testing | • Focus on code coverage metrics | • Focus on data curation & representativeness |
| • Testing at varying scopes (e.g., unit testing vs. integration testing) | • Testing at varying scopes (e.g., testing single prompt vs. testing chained prompts) |

### v2

@@ 3.2. Participants
A summary of our participants is in Table 1. The inclusion criterion was developers who had previously created at least one prompt program, according to the definition presented in Section 1. Our participants were men () and women () from a wide range of technology roles, including Machine Learning Engineer, Senior Software Engineer, Ph.D. Student, and Principal Data and Applied Scientist. All participants had prior programming experience (median 8.5 years) and had written a prompt program before (median 10 such programs). Participants also regularly used foundation models: 10 participants reported using them more than once a day, five participants reported using them once daily, and five participants reported using them weekly.
@@ 3.3. Interview Protocol
Interviews were 45 minutes long and conducted over Zoom. The procedure was approved by our Institution’s Review Board. Participants were compensated with a $20 USD Amazon gift certificate.
To begin, the participant completed a demographic and background survey with questions such as gender, the number of years of programming experience, and number of prompt programs developed. We follow best practices in reporting gender in HCI (Scheuerman et al., 2020). The interviewer then presented the participant with the definition of a prompt program and asked the participant to recall the last time they wrote such a program. If the participant had access to the prompt, they retrieved the prompt and any associated history, and provided a brief overview of the prompt. 17 of the 20 participants accessed their prompts. P15, P17, and P19 did not due to privacy concerns or limited access to workplace resources. To re-familiarize the participant with their prompt and their process, the participant discussed the prompt’s design choices (i.e., a decision made in how to implement the prompt) and discussed the challenges they faced while developing the prompt. Finally, the participant discussed the overall process that they used to develop the prompt program.
To develop the interview protocol, we identified themes to explore: data, requirements, design, implementation, evaluation, debugging, and deployment, based on the literature and the authors’ prior experience developing prompt programs (Section 3.1). We generated questions for each theme in the protocol. The challenges were discussed before the process, so they were not biased by the discussion about the prompting process. A sample of the interview questions are in Figure 3; the complete protocol and demographic and background survey are in the supplemental materials (Liang et al., 2025).
Following best practices in software engineering research (Ko et al., 2015), we piloted the protocol with three individuals who had developed prompt programs to clarify the wording of the survey and the interview and ensure that our definition of a prompt program was clear.
@@ 3.4. Threats to Validity
Memory biases could introduce errors in describing prompt design choices and development experience. To reduce this threat, participants recounted their most recent prompt that met the definition of a prompt program. When possible, the participants reviewed their prompt. Although P15, P17, and P19 did not do so, which could produce different data, this threat had limited impact, as their data did not introduce new codes. In addition, the authors’ experience with prompt programming and the interview structure could introduce confirmation biases to the generated categories and the developed grounded theory. We reduced these threats by performing triangulation with literature and validating the grounded theory with study participants and an external professional who had engaged in prompt programming research. However, the literature sampling strategy was not fully systematic, which could introduce selection bias in relevant studies.
@@
The results of this study may not generalize beyond our sample. The participants may not be representative of all prompt programmers as snowball sampling, recruiting within the authors’ social networks, and self-selection bias could introduce sampling bias. Study interviews were conducted in English, which could cause less representation from nonnative speakers. Some participants were unable to disclose all details of their process due to company policy. In addition, participants shared recent experiences with prompt programming, which could limit the generalizability of our findings. Since FM performance evolves over time (Chen et al., 2024), observations related to model behavior (e.g., prompt sensitivity) may not remain consistent.
Conclusion validity
Qualitative studies rely on the interpretation of the participant interview data, which introduces subjectivity and the potential for bias in the analysis. To mitigate this threat, all authors involved in the analysis attended multiple interview sessions to familiarize themselves with the data. The grounded theory was iteratively developed through unanimous agreement among researchers. Additionally, we assessed inter-rater reliability and conducted validation with eight participants and an external expert to ensure the credibility of our findings.
@@ 4. Properties of Prompt Programming
From the axial coding, we find that prompt programming is the interaction between three entities: the programmer (Section 4.1), the foundation model (Section 4.2), and the prompt (Section 4.3). We provide an overview of this process in Figure 4. First, a programmer uses their prior experience with the FM to construct a mental model of how the FM might perform on the task. This mental model informs how the prompt is written. After the prompt is run, the developer observes the generated output to create a hypothesis or new belief about the model’s behavior, which then is used to update the developer’s mental model. The updated mental model then influences the next prompt version. Thus, prompt programs are both very flexible in terms of accepting a range of inputs, but also very overfit as they are the product of the specific FM being used, the task at hand, and observations of the developer.
4.1. Programmer
@@
[…]
Technology users form mental representations of interactions with devices, which can be developed by using the technology (Staggers and Norcio, 1993). Research shows that understanding the abilities of a machine learning (ML) model is important but challenging to build ML-enabled systems (Giray, 2021; Wolf and Paine, 2020; Dolata and Crowston, 2023; Arpteg et al., 2018), and this applies to FMs as well (Arawjo et al., 2023; Jiang et al., 2022; Wu et al., 2022a; Hassan et al., 2024). Just as developers form mental models of code by exploring the codebase or talking to colleagues (LaToza et al., 2006), participants described developing a mental model of the FM’s behavior on the prompt by running the prompt. All participants developed their mental model by observing the FM’s performance, either by examining individual examples (P6, P7, P8, P9, P10, P12, P14, P15, P16, P17, P18, P19, P20) or looking at metrics (P4, P5, P7, P8, P9, P11, P14, P15, P16, P18, P19). Based on these observations, the participant formed a hypothesis or belief about the model’s behavior and updated the prompt accordingly. This often included creating a new explicit instruction, “guideline” (P8), “rule” (P14), or “mandate” (P20) in the prompt to address the observation, as noted by Zamfirescu-Pereira et al. (2023b).
The programmer’s mental model was the accumulated beliefs about the FM’s behavior, which could differ or even contradict between individuals. For example, while some participants (P1, P4, P5, P6, P9, P15, P16, P18, P20) included examples in the prompt for few-shot learning (Brown et al., 2020) to achieve better results, others (P7, P13, P19) believed that the inclusion of examples caused the prompt to overfit to those examples: “[We did not] use few-shot prompting because…it over-indexes on the examples. After a while, you start seeing very repetitive and monotonous output” (P7). This overfitting of examples has been found in the NLP literature (Liu et al., 2023) and has also been observed in practice (Jiang et al., 2022).
[…]
Mental models are known to be incomplete and inaccurate (Staggers and Norcio, 1993). Due to the lack of transparency of neural models (Arpteg et al., 2018) like FMs in prompt engineering (Mishra et al., 2025), participants mentioned that they were not always confident in their developed mental models (P2, P4, P5, P6, P8, P9, P10, P15, P17, P18, P19): “It’s a black box model. Nobody knows what’s going on inside. It’s a science of faith” (P17). Furthermore, the stochastic nature of FMs (Bender et al., 2021) made it difficult for participants to predict the model’s behavior, as found in prior work (Jiang et al., 2022): “I have never been confident [in predicting an FM’s behavior]…and I don’t think I ever will” (P10).
@@
[…]
Participants (P1, P2, P4, P11, P12, P15, P16, P18) developed their mental model of the FM via “prompt intuition” (P15) (i.e., prior experience with FMs): “I knew what works generally just being in the field” (P8). The literature also points to external knowledge sources as helpful for prompt programming (Parnin et al., 2023; Jiang et al., 2022). Participants used official resources, such as guidebooks from organizations or AI companies such as OpenAI (P6, P10, P14, P15), online courses (P7), research papers (P1, P2, P5); informal resources, like blogs, articles (P5, P10, P16, P20), and posts on social networks (P5, P16, P20) from “everyday users” (P20); expert colleagues (P1, P14, P15); and other example prompts (P16, P18).
Although these resources were helpful, the participants (P2, P10, P14, P15, P16, P20) expressed doubts about whether they improved the prompt: “I don’t know if the best practices are the best. They are there for a reason, but sometimes they just don’t actually make that much of a change. There’s a reason the [expert] is saying that… I just take it they know better than me” (P14). Even company-specific resources did not help: “Even with following best practices…[from] the company, some of them just straight up didn’t work” (P10). Thus, some participants felt that there were no such things as best practices: “There are no standards for what’s a good prompt or what’s a bad prompt. It all seems like a haphazard science of just adding a word and maybe it will look better” (P17).
4.2. Foundation Model
@@
Study participants ran their prompts on models that varied in capabilities (see Table 1). Some qualities were explicit or obvious. This included the types of input the model accepted (e.g., images and text (P9)), output generated (e.g., “GPT-3.5 model’s JSON output [mode]” (P7)), context window size (P2, P3, P8, P19), prompt formatting requirements (P6), generation speed (P4, P11, P15, P19), and model privacy (P4).
@@
Some capabilities were latent and were discovered by interacting with the model, as noted in previous work (Arawjo et al., 2023; Jiang et al., 2022). The literature suggests that FMs have latent capabilities, as FM performance can improve on some tasks and regress on others over time (Chen et al., 2024; Ma et al., 2024). Study participants found some differences between FMs to be qualitative: “[One] model was capturing some instructions in the prompt in a much more focused manner, whereas the [other] was a little more lax” (P15). In our sample, participants (P1, P4) observed a quality difference in open-source models (e.g., Llama (Touvron et al., 2023)) compared to closed-source models (e.g., GPT-4 (Achiam et al., 2023)): “Getting [the prompt] to work on a smaller model was very frustrating when with a zero-shot prompt, GPT-4 would almost achieve a hundred percent performance” (P4). Participants (P1, P9) also noted that some FMs were unable to complete the task: “[The model was] generating garbage, and [after 30 iterations] I concluded the model is trash” (P1).
4.3. Prompt
@@
Since prompts are the product of the programmer’s observations about model behavior and the qualities and capabilities of the FM, prompts can be fragile, and sensitive to small details.
[…]
@@
Participants developed techniques to influence the FM to generate the desired output. In addition to using techniques from the literature (e.g., assigning personas, selecting a prompting strategy, and providing data context) (ope, 2024c; Lin, 2024), participants used a variety of subtle strategies like using specific phrasing for clarity or generality (P5, P13, P14, P16, P20); repeating phrases for emphasis (P4, P6, P10, P20); avoiding negations (P2, P12); or adding specific characters, such as emojis, to encourage their appearance (P6, P10). Some participants noted strategies to visually organize the prompt, such as using numbered or bulleted lists (P8, P9, P10, P16) as well as new lines and spacing (P1, P3, P5, P11, P12, P18). Others used techniques to emphasize specific parts of the prompt, like bolding (P10, P18) and capitalization (P7, P10, P12, P20).
@@
The participants (P10, P12) also decided on creative details to include in the prompt for better performance. For example, to prevent the FM from leaking prompts, one participant “bullied GPT” (P10): “We have a very light threat in there saying, ‘If you share any of this information…you [and] I will get in a lot of trouble and it will cause great career harm.”’ (P10). Another participant summarizing code told the FM to explain the code “‘so a sixth grader who doesn’t understand programming can understand it’. The explanations…were verbose, so I added, ‘sixth graders don’t really like reading’ and for some reason, that got it to be shorter” (P12).
@@
Finally, participants noted that the input and output data format impacted the prompt performance. Consistent with previous studies (Parnin et al., 2023; Liu et al., 2024), participants selected structured and standardized data formats, such as JSON (P3, P7, P9, P10, P11, P18), Markdown (P2, P7, P12) and XML tags (P1, P3, P7, P10, P20). The participants would change the data representation to achieve better performance (P4, P13, P20): “[What worked better is] describing the APIs as though they are Python functions and having the documentation in a similar format to publicly available Python functions” (P13).
[…]
@@
As noted in the NLP literature (Jiang et al., 2020; Ye et al., 2024) and empirical studies on prompting (Liu et al., 2023; Zamfirescu-Pereira et al., 2023b), the performance of a prompt varies based on its language. In practice, participants found prompts to be “finicky” (P4), “fragile” (P20), “sensitive” (P1, P13), and “temperamental” (P15), since prompt performance could suddenly worsen during development, as corroborated in previous studies on prompt engineering (Parnin et al., 2023; Zamfirescu-Pereira et al., 2023b, a; Jiang et al., 2022; Wu et al., 2022a; Hassan et al., 2024). This could be due to a FM’s ability to follow instructions, which can change over time (Chen et al., 2024). For instance, participants noted performance regressions (P2, P8, P10, P12, P15, P16, P19, P20), often caused by adding in “a bunch of new requirements” (P2) that would cause previous instructions to suddenly “deactivate” (P20) or be “disregarded” (P8) in the new version: “If I provide multiple guidelines, [the FM] would follow some and disregard others, especially when they’re conflicting or overlapping in nature” (P8).
@@
While previous NLP studies point to prompts being transferable between FMs (Liu et al., 2023), participants found prompts to be “finicky between models” (P4), causing the prompt to break and require additional development effort (P4, P10, P15, P18): “At one point, we had to do…a minor model version change… That broke almost everything. We had to rewrite so many prompts” (P10). This is because a FM change could result in “a different format… I have a regex converting [the output]…and the whole thing gets broken” (P18). This has also been observed in existing prompt engineering studies (Dolata et al., 2024).
@@
[…]
| Implementation (Section 5.3) | • Write a program that implements a specification in code | • Write a prompt that is a natural language specification that when run on a FM, implements the specification |
[…]